Complete the following reactions and balance them:

1. B + H2So4 (concentrated )-----

2. LiBH4 + H2o ---------
1
Expert's answer
2019-08-12T07:39:47-0400

1) Boron will be oxidized with concentrated sulfuric acid up to boric acid formation:

B + H2SO4 = H3BO3 + SO2 + H2O

SO42- + 4H+ + 2e = SO2 + 2H2O

B + 3H2O - 3e = BO33- + 6H+

Thus,

2B + 3H2SO4 = 3SO2 + 2H3BO3

2) Lithium tetrahydroborate will be completely hydrolised and the products will react to form lithium borate and hydrogen gas:

LiBH4 + 2H2O = LiBO2 + 4H2